OpenStudy (anonymous):

someone please help solve simultaneously: 7/x - 5/y=3 and 2/x + 25/2y=12

OpenStudy (anonymous):

it is easy consider u=1/x and v=1/y, then the equation will become 7u-5v=3 4u+25v=24, then multiply the equation one with 5 yu will get 35u-25v=15 and add the last two you will get 39u=39 then u=1 and 35-15=25v i.e. 20=25v so v=20/25=4/5 then x=1/u, which gives x=1 and y=1/v which is y=5/4

OpenStudy (anonymous):

2x-2y=0